OpenStudy (anonymous):

Find the sum of trigonometric equation solution in the range of [0,2pi]:

OpenStudy (anonymous):

\[\cos2x+\sin^2x=\frac{ 3 }{ 4 }\]

OpenStudy (anonymous):

@AravindG @ganeshie8 @mathstudent55 @eliassaab @thomaster @SolomonZelman @hoblos

OpenStudy (anonymous):

\[cosx= \frac{ \sqrt{3} }{ 2 }\]

OpenStudy (anonymous):

but i don't have the full answer

Parth (parthkohli):

\[\cos(2x) = \cos^2 (x) - \sin^2(x)\]

OpenStudy (anonymous):

i solved the exercise but i don't know how to finalize it

Parth (parthkohli):

OK. Do you know the angles for which \(\cos(x) = \dfrac{\sqrt{3}}{2}\)?

OpenStudy (anonymous):

yes pi/6 i guess..

Parth (parthkohli):

Good job.

OpenStudy (anonymous):

actually it's not

OpenStudy (anonymous):

Read again the exercise!

Parth (parthkohli):

hmm, isn't that what the question asks for?

Parth (parthkohli):

Oh, it looks for the sum of all angles. I think.

OpenStudy (anonymous):

yes

Parth (parthkohli):

OK, then look at your unit circle.

OpenStudy (anonymous):



Parth (parthkohli):

\[\cos(2\pi - x) = \cos(x)\]So \(2\pi - \pi/6\) also is an angle.

OpenStudy (anonymous):

so do you mean cos it's a periodical function of 2pi period?

Parth (parthkohli):

OK, let's look at the unit circle. That'd explain everything. cos is the x-coordinate on the unit circle, right?

OpenStudy (anonymous):

right

Parth (parthkohli):